My problem is this:
Given the numbers 4, -7, 3/4, pie, -0.2, list those that are rational. I'm assuming they are all rational except for pie. As 4 can be written as 4/1 etc.
thanks.

Answer by tutorcecilia(2144)

I agree because pi cannot be converted to a repeating or terminating decimal; nor can pi be converted to a fraction.
.
All the other numbers are rational because:
4 = 4/1 = 4.00
-7 = -7/1 = -7.00
3/4 = .25
-0.2 = -1/5